# Towards Training-Free Open-World Classification with 3D Generative Models

### v1

@@
Abstract
3D open-world classification is a challenging yet essential task in dynamic and unstructured real-world scenarios, requiring both open-category and open-pose recognition. To address these challenges, recent wisdom often takes sophisticated 2D pre-trained models to provide enriched and stable representations. However, these methods largely rely on how 3D objects can be projected into 2D space, which is unfortunately not well solved, and thus significantly limits their performance. Unlike these present efforts, in this paper we make a pioneering exploration of 3D generative models for 3D open-world classification. Drawing on abundant prior knowledge from 3D generative models, we additionally craft a rotation-invariant feature extractor. This innovative synergy endows our pipeline with the advantages of being training-free, open-category, and pose-invariant, thus well suited to 3D open-world classification.
Extensive experiments on benchmark datasets demonstrate the potential of generative models in 3D open-world classification, achieving state-of-the-art performance on ModelNet10 and McGill with 32.0% and 8.7% overall accuracy improvement, respectively. The code is available in the supplementary materials.
1 Introduction
Figure 1: Comparison of Pipelines: (a) SOTA methods project 3D samples into 2D images to harness 2D prior knowledge for open category classification but are sensitive to pose changes, (b) The proposed pipeline applies 3D prior knowledge to generate anchor samples for novel categories, embedding them in a rotation-invariant space for effective performance in open-pose scenarios.
@@ 1 Introduction
3D classification within open-world environments is essential for developing computational models that can accurately interpret complex and unstructured real-world scenarios. Such advances are crucial for improving the autonomy and adaptability of intelligent systems Liu et al. (2023); Zhang et al. (2023); Zhao et al. (2024). This field aims to effectively recognize objects facing various real-world challenges, particularly those involving unexpected novel categories (open-category) and arbitrary object orientations (open-pose) Zhang et al. (2023); Zhu et al. (2023); Zhao et al. (2025).
@@
Current efforts in open-world 3D classification rely primarily on transfer learning methods, categorized into language-only, language-point, and language-image model-based strategies Zhao et al. (2025). While these strategies showcase innovative methodologies, each encounters specific hurdles that inspire our investigation. Language-only model-based methods Cheraghian et al. (2020, 2022) rely solely on textual information to infer connections with novel classes, often fall short due to the limited descriptive power of text and the oversight of essential spatial information inherent in 3D data. Language-point model-based methods Xue et al. (2023, 2024); Xu et al. (2024), on the other hand, demands extensive data and significant computational resources for training, posing practical limitations. Language-image model-based methods Zhu et al. (2023); Huang et al. (2023); Hegde et al. (2023); Shen et al. (2024); Zhao et al. (2025) further complicate matters by requiring intricate projection hyperparameters to effectively map 3D data to 2D images. In particular, all existing approaches exhibit sensitivity to pose changes.
Recognizing these issues, we argue that the field currently faces a pressing need to effectively harness the spatial information inherent in 3D data while maintaining computational efficiency. In response to this challenge, we explore leveraging the rich prior knowledge of advanced 3D generative models Luo and Hu (2021); Nichol et al. (2022); Jun and Nichol (2023); Yi et al. (2024) by introducing an innovative pipeline that formulates a succinct and efficient 3D classifier for open-world tasks. This pipeline utilizes pre-trained 3D generative and representation models for object classification to achieve efficient performance without additional training, significantly reducing computational demands and data requirements. As shown in Fig. 1, the state-of-the-art (SOTA) methods employ language-image models, which project 3D samples into 2D images to exploit 2D prior knowledge for the classification of open categories. However, the generalizability of the 2D knowledge limits recognition, especially in open-pose scenarios where the object’s orientation and the projection style can affect the outcomes. Conversely, our proposed pipeline capitalizes on 3D prior knowledge to generate anchor samples for novel categories, subsequently embedding them into a rotation-invariant space to mitigate the 3D information loss inherent in projection settings.
@@
The execution of our pipeline begins with feeding textual descriptions into a pre-trained text-to-3D generation model, which synthesizes anchor samples serving as prototypical examples for each category. Subsequently, a pre-trained point cloud representation model is utilized to extract anchor features from these samples. These anchor features are indicative of the classification capacity of our developed classifier, as its performance exhibits a positive correlation with the quality and diversity of these generated samples (see Section  4.3). Ideally, a sufficiently generalized generator will enhance the proficiency of our classifier in accurately categorizing any 3D objects. During the inference phase, cosine similarity within the representation space is computed to evaluate the similarity between test samples and anchor samples, thus leading to category predictions for the test samples. Further, we investigate the impact of various representation models on open-world classification. Our findings suggest that most representation models fail to provide consistent representations for identical 3D objects across different orientations, displaying rotational variance. Extensive experiments on the McGill and ModelNet10 datasets validate the robustness of our approach. Additionally, we explore the use of large language models (LLMs) to generate detailed textual descriptions of category names, with the aim of enhancing limited textual descriptive capabilities.
@@
  • Utilizing 3D point cloud augmentation techniques, we have achieved state-of-the-art performance on ModelNet10 and McGill with overall accuracy improvement of 32.0% and 8.7%, respectively.
@@
Figure 2: Overview of the network architecture with two components. (1) Open-world classifier: Category names are input into ChatGPT to create descriptions, utilized by a pre-trained text-to-3D model to generate anchor samples. These samples are augmented to align with test samples. A pre-trained 3D model extracts features, marking them as prototypes for classification. (2) Test sample inference: Test samples are augmented and processed through the same model for feature extraction. Predictions are based on cosine similarity between test and anchor sample features. Different colors denote new categories, gray represents test data.
2 Related Work
2.1 3D Open-World Classification
@@ 2 Related Work
Current 3D open-world classification methods can be divided into three categories Zhao et al. (2025). Language-only model-based methods Cheraghian et al. (2020, 2022); Hao et al. (2023) utilize pre-trained language models like Word2Vec Mikolov et al. (2013) or GloVe Pennington et al. (2014) to establish semantic connections between training and novel classes but have fallen out of favor due to reliance on a single source of prior knowledge. Language-image model-based methods Zhu et al. (2023); Huang et al. (2023); Hegde et al. (2023); Shen et al. (2024); Zhao et al. (2025) project point clouds into multiple 2D views and utilize pre-trained models like CLIP Radford et al. (2021) or Diffusion Rombach et al. (2022); Sohl-Dickstein et al. (2015) for classifying these projections, enabling 3D open-world classification. Although effective, they involve numerous hyperparameters, such as projection angles and viewpoint weights, which complicate the process and diminish robustness. Language-point model-based methods Xue et al. (2023, 2024); Xu et al. (2024) link point clouds with text through various datasets and large language models. These methods need significant data resources and GPU memory, often with 7 or 13 billion parameter models. A recent study Zhao et al. (2025) highlights the limitations of existing methods in handling open-pose 3D objects in real-world applications. Conversely, our research initiates an innovative approach within the framework of language-point model methodologies by investigating current text-to-3D generative models Nichol et al. (2022); Jun and Nichol (2023); Yi et al. (2024) to address 3D open-world classification. Benefiting from the prior knowledge inherent in 3D generative models, our method achieves SOTA performance in open-pose scenarios without extensive hyperparameter tuning or large datasets for training.
2.2 Text-to-3D Generation
@@
With the success of diffusion models in image generation Nichol et al. (2021), studies Luo and Hu (2021); Nichol et al. (2022); Jun and Nichol (2023); Yi et al. (2024) now integrate them into text-to-point-cloud tasks. Some methods Nichol et al. (2022); Jun and Nichol (2023); Yi et al. (2024) follow a ”text-to-image-to-point-cloud” paradigm, wherein a 2D image is first generated through the diffusion model and then converted into a 3D point cloud. Alternatively, some approaches Luo and Hu (2021); Ho et al. (2020) synthesize point clouds directly in 3D by sampling points and using a diffusion model. These pre-trained 3D generative models, rich in prior knowledge, can potentially serve as classifiers in open-world scenarios. We validate this using Shap-e Jun and Nichol (2023) and GaussianDreamer Yi et al. (2024), chosen for their speed and performance.
2.3 3D Represent Learning
@@
Initial approaches to feature extraction from point clouds relied on geometric methods utilizing local information like normals and curvature Rister et al. (2017). PointNet Qi et al. (2017) was the first to use neural networks for spatial feature extraction. Inspired by the Transformer’s success in image and text domains Vaswani et al. (2017), many studies Wang (2023); Wu et al. (2024); Mehraban et al. (2024); Jiang et al. (2025) adopted its framework for 3D point cloud models, leveraging its attention mechanism to excel in capturing both local and global features. Addressing the open-pose challenge in open-world scenarios Zhao et al. (2025), we compare two models: TAP Wang et al. (2023), which lacks rotation invariance, and TET Melnyk et al. (2024), which tackles it. TET’s core idea involves a nonlinear TTlayer ($l_{TT}$), mapping point clouds from $\mathbb{R}^{N\times 3}$ to $\mathbb{R}^{N\times 4\times K}$ via steerable spherical neurons, embedding data into conformal space to ensure rotational and permutation equivariance.
3 Method
3.1 Problem Statement
@@
Given $N_{c}$ categories with the names set $\mathbf{CN}=\{CN_{1},CN_{2},...,CN_{N_{c}}\}$, the 3D open-world classification task involves assigning the correct category label in $\mathbf{CN}$ to each instance within a sample set $\mathbf{TS}\in\mathbb{R}^{N_{t}\times N_{p}\times 3}$. Here, $N_{t}$ refers to the number of point cloud test samples, and $N_{p}$ implies the number of points in each sample. Typically, in an open-world setting, we are unable to obtain real samples corresponding to these categories to train a classifier.
3.2 Overall
The overall network architecture of the proposed method is depicted in Fig. 2. It consists primarily of two components: the construction of an open-world classifier and inferring test samples. Initially, we provide the given category text names as input to ChatGPT to retrieve comprehensive textual descriptions of each category. Subsequently, these textual descriptions are input into a pre-trained text-to-3D model to generate anchor samples corresponding to the respective categories. Data augmentation techniques are employed to process the anchor samples, thereby reducing the geometric representation differences between the anchor and test samples. Furthermore, we utilize the pre-trained 3D representation model to extract feature representations from these anchor samples. These extracted features serve as the prototypical examples for our open-world classifier. During the inference phase, we apply the same data augmentation procedures to the inference samples and input the augmented samples into the same representation model to obtain their feature representations. By comparing the cosine similarity between the feature representations of the test samples and the anchor samples, we can obtain category predictions for the test samples. More details follow in later subsections.
3.3 Construct the 3D Open-World Classifier
@@
$\mathbf{CD}=\mathbf{LLMs}\left(Q,\mathbf{CN}\right),$ (1)
@@
$AS_{i}^{j}=\mathbf{GMs}\left(noise_{j},CD_{i}\right),\quad AS_{i}^{j}\in\mathbb{R}^{N_{p}^{{}^{\prime}}\times 3},$ (2)
@@
$\hat{AS_{i}^{j}}=\frac{\mathbf{FPS}(AS_{i}^{j},\hat{N_{p}})-\mathbf{Avg}(AS_{i}^{j})}{\mathbf{L}(AS_{i}^{j})}\cdot R,\quad R\in\mathbb{R}^{3\times 3},$ (3)
@@
Anchor Sample Feature Distribution. Different 3D representation models $\mathbf{RMs}$ have a significant impact on 3D open-world classifiers. Existing models are often limited to aligned datasets and overlook the influence of open-pose variations. After thorough exploration, we select TET Melnyk et al. (2024), a 3D representation model that focuses on the study of rotational variance, as the feature extractor for constructing our open-world classifier. To this end, the anchor sample feature $AF$ could be attained by the following formula:
$AF_{i}^{j}=\mathbf{RMs}\left(\hat{AS_{i}^{j}}\right),\quad AF_{i}^{j}\in\mathbb{R}^{\hat{N_{p}}\times D},$ (4)
where $AF_{i}^{j}$ serves as the $j_{th},j\in(0,N_{a}]$ anchor sample feature for the $i_{th}$ category, and $D$ donates the feature dimension. As shown in Fig. 2, the anchor sample feature distribution implies the classification ability of our 3D open-world classifier. The hyperplane separating the features acts as the decision boundary for our classifier. Furthermore, the region where the features of a test sample fall determines the predicted category of that test sample. In the following section, we offer a comprehensive explanation of how our constructed open-world classifier is employed to classify objects.
3.4 Inference Phase

### v2

@@
Abstract.
3D open-world classification is a challenging yet essential task in dynamic and unstructured real-world scenarios, requiring robust subsequent knowledge adaptation capabilities. While current approaches predominantly rely on 2D pre-trained models through 3D-to-2D projection, their performance degrades severely under arbitrary object orientations. Unlike these present efforts, this work makes a pioneering exploration of 3D generative models for 3D open-world classification-specifically, leverageing the accumulated prior knowledge from these models to provide anchors for novel categories, while integrating a rotation-invariant feature extractor. This innovative synergy endows our pipeline with the advantages of being training-free and pose-invariant, thus well suited to adapt novel categories in 3D open-world classification. Extensive experiments on benchmark datasets demonstrate the potential of this pipeline, achieving state-of-the-art performance on ModelNet10 and McGill with 32.7% and 8.7% overall accuracy improvement, respectively. The code is available in the supplementary materials.
3D Open-World Classification, 3D Generative Models, Training-Free Framework, Pose-Invariant Recognition
ccs: Computing methodologies Object recognition
1. Introduction
@@
3D classification in open-world environments presents critical challenges in developing computational models that can accurately interpret complex, unstructured scenarios while adapting to novel knowledge. Such advances are crucial for improving the autonomy and adaptability of intelligent systems (Liu et al., 2023; Zhang et al., 2023; Zhao et al., 2024). This field aims to effectively categorize objects that not only vary in trained and novel categories–a challenge of open-category recognition–but also appear in arbitrary poses, posing an open-pose recognition dilemma (Zhang et al., 2023; Zhu et al., 2023; Zhao et al., 2025).
@@
Knowledge adaptation in current open-world 3D classification approaches is based mainly on three paradigms: language-only, language-point, and language-image model-based strategies (Zhao et al., 2025). While promising, each suffers from limitations that motivate our investigation. Language-only model-based methods (Cheraghian et al., 2020, 2022) rely solely on textual information to infer connections with novel classes, often fall short due to the limited descriptive power of text and the oversight of essential spatial information inherent in 3D data. Language-point model-based methods (Xue et al., 2023; Xue et al., 2024; Xu et al., 2024), on the other hand, demands extensive data and significant computational resources for training, posing practical limitations. Language-image model-based methods (Zhu et al., 2023; Huang et al., 2023; Hegde et al., 2023; Shen et al., 2024; Zhao et al., 2025) further complicate matters by requiring intricate projection hyperparameters to effectively map 3D data to 2D images. In particular, all existing approaches exhibit sensitivity to pose changes.
Figure 1. Comparison of Pipelines: (a) SOTA methods (Zhao et al., 2025; Zhu et al., 2023) project 3D samples into 2D images to harness 2D prior knowledge for novel category recognition but are sensitive to pose changes, (b) The proposed pipeline applies 3D prior knowledge to generate anchor samples for novel categories, embedding them in a rotation-invariant space for effective performance in open-pose scenarios. Predictions are computed via feature similarity between test and anchor samples.
Addressing these limitations requires meeting two critical demands: 1) effectively exploiting the spatial information in 3D data while maintaining computational efficiency, and 2) achieving strong adaptation capability for novel knowledge in dynamic environments. Current approaches typically require extensive retraining when encountering new scenarios - a significant drawback in open-world settings where environmental factors (like object orientation) constantly evolve.
In response to these challenges, we present a novel training-free framework that synergistically leverages pre-trained 3D generative models (Luo and Hu, 2021; Nichol et al., 2022; Jun and Nichol, 2023; Yi et al., 2024) with representation learning to construct a lightweight yet efficient 3D classifier for open-world scenarios. Our proposed pipeline capitalizes on 3D prior knowledge to generate anchor samples for novel categories, subsequently embedding them into a rotation-invariant space. As shown in Fig. 1, unlike the state-of-the-art (SOTA) methods that rely on projecting 3D data into 2D space to exploit limited 2D priors from vision-language models, our method mitigates two critical limitations: 1) the inherent information loss from 3D to 2D projection that particularly handicaps pose-variant recognition, and 2) the computational overhead of continuous model retraining for new scenarios.
@@
The execution of our pipeline begins with feeding textual descriptions into a pre-trained text-to-3D generation model, which synthesizes anchor samples serving as prototypes for each category. Subsequently, a pre-trained point cloud representation model is utilized to extract anchor features from these samples. These anchor features are indicative of the classification capacity of our developed classifier, as its performance exhibits a positive correlation with the quality and diversity of these generated samples (see Section  4.3). Ideally, a sufficiently generalized generator will enhance the proficiency of our classifier in accurately categorizing any 3D objects. During the inference phase, cosine similarity within the representation space is computed to evaluate the similarity between test samples and anchor samples, thus leading to category predictions for the test samples. In addition, we investigate the impact of various representation models on open-world classification. Our findings suggest that most representation models fail to provide consistent representations for identical 3D objects across different orientations, displaying rotational variance. Extensive experiments on the McGill and ModelNet10 datasets validate the robustness of our approach. Additionally, we explore the use of large language models (LLMs) to generate detailed textual descriptions of category names, with the aim of enhancing limited textual descriptive capabilities.
@@
  • Utilizing 3D point cloud augmentation techniques, we achieve state-of-the-art performance on ModelNet10 and McGill with overall accuracy improvement of 32.7% and 8.7%, respectively.
@@
Figure 2. Overview of the network architecture with two components. (1) Open-world classifier: Category names are input into ChatGPT to create descriptions, utilized by a pre-trained text-to-3D model to generate anchor samples. These samples are augmented to align with test samples. A pre-trained 3D model extracts features, marking them as prototypes for classification. (2) Test sample inference: Test samples are augmented and processed through the same model for feature extraction. Predictions are based on cosine similarity between test and anchor sample features. Different colors denote new categories, gray represents test data.
2. Related Work
2.1. 3D Open-World Classification
@@ 2. Related Work
Current 3D open-world classification methods can be divided into three categories (Zhao et al., 2025). Language-only model-based methods (Cheraghian et al., 2020, 2022; Hao et al., 2023) utilize pre-trained language models like Word2Vec (Mikolov et al., 2013) or GloVe (Pennington et al., 2014) to establish semantic connections between training and novel classes but have fallen out of favor due to reliance on a single source of prior knowledge. Language-image model-based methods (Zhu et al., 2023; Huang et al., 2023; Hegde et al., 2023; Shen et al., 2024; Zhao et al., 2025) project point clouds into multiple 2D views and utilize pre-trained models like CLIP (Radford et al., 2021) or Diffusion (Rombach et al., 2022; Sohl-Dickstein et al., 2015) for classifying these projections, enabling 3D open-world classification. Although effective, they involve numerous hyperparameters, such as projection angles and viewpoint weights, which complicate the process and diminish robustness. Language-point model-based methods (Xue et al., 2023; Xue et al., 2024; Xu et al., 2024) link point clouds with text through various datasets and large language models. These methods need significant data resources and GPU memory, often with 7 or 13 billion parameter models. A recent study (Zhao et al., 2025) highlights the limitations of existing methods in handling open-pose 3D objects in real-world applications. Conversely, our research initiates an innovative approach within the framework of language-point model methodologies by investigating current text-to-3D generative models (Nichol et al., 2022; Jun and Nichol, 2023; Yi et al., 2024) to address 3D open-world classification. Benefiting from the prior knowledge inherent in 3D generative models, our method achieves SOTA performance in open-pose scenarios without extensive hyperparameter tuning or large datasets for training.
2.2. Text-to-3D Generation
@@
With the success of diffusion models in image generation (Nichol et al., 2021), studies (Luo and Hu, 2021; Nichol et al., 2022; Jun and Nichol, 2023; Yi et al., 2024) now integrate them into text-to-point-cloud tasks. As the pioneering work, Diffusion-PC (Luo and Hu, 2021) introduces a novel probabilistic framework for generating 3D point clouds. By modeling the reverse diffusion process as a Markov chain, inspired by non-equilibrium thermodynamic principles, it effectively transforms noise distributions into structured point clouds. Some methods (Nichol et al., 2022; Jun and Nichol, 2023; Yi et al., 2024) follow a ”text-to-image-to-point-cloud” paradigm, wherein a 2D image is first generated through the diffusion model and then converted into a 3D point cloud. For instance, in STPD (Wu et al., 2023), a sketch and text guided diffusion model for colored point cloud generation is proposed, which first generates a 2D image and then converts it into a 3D point cloud. This approach leverages the strengths of diffusion models in image generation and then uses additional techniques to transform the 2D image into a 3D point cloud. Alternatively, some approaches (Luo and Hu, 2021; Ho et al., 2020) synthesize point clouds directly in 3D by sampling points and using a diffusion model. These pre-trained 3D generative models, rich in prior knowledge, can potentially serve as classifiers in open-world scenarios. In this work, we validate this using Shap-e (Jun and Nichol, 2023) and GaussianDreamer (Yi et al., 2024), chosen for their speed and performance. Specifically, Shap-e (Jun and Nichol, 2023) is known for its ability to directly generate point clouds based on given prompts, while GaussianDreamer (Yi et al., 2024) first generates a coarse point cloud and then optimizes it using 3D Gaussian Splatting. Both models could be used to generate point clouds that can serve as inputs for classification tasks in open-world scenarios.
2.3. 3D Represent Learning
@@
Initial approaches to feature extraction from point clouds relied on geometric methods utilizing local information like normals and curvature (Rister et al., 2017). PointNet (Qi et al., 2017) is the first to use neural networks for spatial feature extraction. Inspired by the Transformer’s success in image and text domains (Vaswani et al., 2017), many studies (Wang, 2023a; Wu et al., 2024; Mehraban et al., 2024; Jiang et al., 2025) adopt its framework for 3D point cloud models, leveraging its attention mechanism to excel in capturing both local and global features. For instance, Point Transformer (Zhao et al., 2021) is designed to take advantage of the local geometric relations between the center point and its neighbors. Moreover, multi-view (Hu et al., 2021; Yang et al., 2023), voxel (Maturana and Scherer, 2015; Choy et al., 2019; Zhao et al., 2023), and octree (Wang et al., 2017; Wang, 2023b) representations are also popular forms of point cloud representation. Addressing the open-pose challenge in open-world scenarios (Zhao et al., 2025), we compare two models: TAP (Wang et al., 2023), which lacks rotation invariance, and TET (Melnyk et al., 2024), which tackles it. TAP (Wang et al., 2023) is a novel 3D-to-2D generative pre-training method for point cloud models. TAP generates 2D view images from 3D point clouds using a cross-attention mechanism, significantly improving performance on downstream tasks like classification and segmentation. Moreover, TET’s (Melnyk et al., 2024) core idea involves a nonlinear TTlayer ($l_{TT}$), mapping point clouds from $\mathbb{R}^{N\times 3}$ to $\mathbb{R}^{N\times 4\times K}$ via steerable spherical neurons, embedding data into conformal space to ensure rotational and permutation equivariance.
3. Methodology
3.1. Problem Statement
@@
Given $N_{c}$ categories with the names set $\mathbf{CN}=\{CN_{1},CN_{2},...,\\ CN_{N_{c}}\}$ in a dynamic environment, the 3D open-world classification task involves assigning the correct category label in $\mathbf{CN}$ to each instance within a sample set $\mathbf{TS}\in\mathbb{R}^{N_{t}\times N_{p}\times 3}$. Here, $N_{t}$ and $N_{p}$ refers to the number of point cloud test samples and the number of points in each sample. Typically, in open-world settings, real samples for these categories are unavailable for classifier training. To bridge this gap, external knowledge can be leveraged to support adaptation to these novel categories.
3.2. Overall Framework
As illustrated in Fig. 2, our framework introduces a training-free adaptation paradigm for open-world 3D classification comprising two components: the construction of an open-world classifier and inferring test samples. This paradigm uniquely addresses environmental dynamics through three key mechanisms: 1) Knowledge-Aware Anchor Synthesis: We first leverage large language models (e.g., ChatGPT) to obtain rich semantic descriptions for each novel category name, which are then transformed into diverse 3D anchor samples using pre-trained text-to-3D generative models. Crucially, this process preserves the full 3D spatial information that would be lost in projection-based approaches. 2) Dynamic Environment Adaptation: To robustly handle pose variations and environmental dynamics, we employ a carefully designed augmentation strategy that generates geometrically transformed variants of each anchor sample. These augmented samples are then processed through a frozen pre-trained 3D encoder to obtain rotation-invariant feature prototypes for each category. 3) Efficient Open-World Inference: For each test sample, we compute its feature representation and determine its category via cosine similarity to our prototypical anchors. This lightweight comparison mechanism eliminates the need for costly retraining while maintaining adaptation capability to novel environmental conditions.
The entire pipeline operates without retraining, making it particularly suitable for dynamic environments where new categories and viewing conditions may emerge frequently. As shown in Tab. 1, this approach achieves superior performance. Fig. 3 demonstrates its high efficiency - this is achieved with only 7 samples per class, significantly fewer than required by retraining approaches.
3.3. Construct the 3D Open-World Classifier
@@
(1) $\mathbf{CD}=\mathbf{LLMs}\left(Q,\mathbf{CN}\right),$
@@
(2) $AS_{i}^{j}=\mathbf{GMs}\left(noise_{j},CD_{i}\right),\quad AS_{i}^{j}\in\mathbb{R}^{N_{p}^{{}^{\prime}}\times 3},$
@@
(3) $\hat{AS_{i}^{j}}=\frac{\mathbf{FPS}(AS_{i}^{j},\hat{N_{p}})-\mathbf{Avg}(AS_{i}^{j})}{\mathbf{L}(AS_{i}^{j})}\cdot R,\quad R\in\mathbb{R}^{3\times 3},$
@@
Anchor Sample Feature Distribution. Different 3D representation models $\mathbf{RMs}$ have a significant impact on 3D open-world classifiers. Existing models are often limited to aligned datasets and overlook the influence of open-pose variations. After thorough exploration, we select TET (Melnyk et al., 2024), a 3D representation model that focuses on the study of rotational variance, as the feature extractor for constructing our open-world classifier. To this end, the anchor sample feature $AF$ could be attained by the following formula:
(4) $AF_{i}^{j}=\mathbf{RMs}\left(\hat{AS_{i}^{j}}\right),\quad AF_{i}^{j}\in\mathbb{R}^{\hat{N_{p}}\times D},$
where $AF_{i}^{j}$ serves as the $j_{th},j\in(0,N_{a}]$ anchor sample feature for the $i_{th}$ category, and $D$ donates the feature dimension. As shown in Fig. 2, the anchor sample feature distribution implies the classification ability of our 3D open-world classifier. The hyperplane separating the features acts as the decision boundary for our classifier. On the other hand, the region where the features of a test sample fall determines the predicted category of that test sample. In the following section, we offer a comprehensive explanation of how our constructed open-world classifier is employed to classify objects.
3.4. Inference Phase